# Literature Review: Cyber Security Monitoring in Maritime

### v2

@@ V-D Security monitoring methods for maritime CPS
Pelissero et al. [49] created and utilised graph-based models for the detection of attack propagation in CPSs. A graph model that represented the digital and physical system layers was proposed in the paper, considering the interdependencies between CPSs by using the system variables. The model was applied in a case study that involved a small-scale system consisting of a ship with a fuel tank, valve, motor and propeller which were monitored through a PLC. One scenario of normal system usage and four attack scenarios including port scan, closing the PLC and man-in-the-middle attacks were simulated in the study. According to the authors, the method did not produce a false positive alert for the normal scenario, whereas one attack scenario out of four remained undetected (i.e., precision and recall were 100% and 75%, respectively). However, the evaluation of the method was limited to the aforementioned five scenarios without testing the method on a larger dataset.
@@
The study by Boudehenn et al. [48] focused on the use of Teager-Kaiser operator for time series analysis to detect anomalies in a ship’s propulsion system in the case of cyber attacks. During the experiments, the authors collected the following time series data – oil consumption, fuel consumption, propeller speed, and valve opening. The following attacks were conducted for evaluating the anomaly detection methods – port scanning with nmap tool, DoS attack with hping tool, and exploiting the vulnerability of the ship’s propulsion PLC for stopping and starting the PLC. Port scanning and DoS attacks did not influence the propulsion system, and only the attack against the PLC had a significant impact. According to the experiments, the Teager-Kaiser operator managed to identify sudden changes in the ship’s propulsion system time series data. However, the evaluation results were presented in the form of graphs only for the 5-minute time period which involved a single attack against the PLC. Therefore, it is difficult to assess what would be the detection rate for many attacks of different types against the propulsion system during a longer time frame (e.g., several days or weeks), and a more detailed analysis of the proposed method was identified as future work in the study.
The study by Qiu et al. [46] proposed the use of a Particle filter-based anomaly detection method for time series data collected from a ship’s CPS to detect cyber attacks. In the study, the authors focused on data collected from the ship’s propulsion system – the temperature of the cylinder exhaust, cooling water pressure, and the speed of the fuel flow. For training the anomaly detection system, the authors used data collected from a new cargo ship during its first sailing week, assuming that the probability of cyber attacks during this week is very low. For evaluating the anomaly detection algorithm, data from the second half of the first sailing year was employed. For discussing the anomaly detection rate of the algorithm, the authors presented graphs for the 5,000-minute time frame (about 3.5 days) in the study. Although the graphs displayed some fluctuations in the time series data for the propulsion system, the study did not detail whether these sudden changes were the result of a cyber attack or rather a side effect of normal navigation.
@@ V-E Security monitoring system architectures
The study by Jacq et al. [13] proposed a maritime SOC architecture. The authors first provided an overview of the specific constraints of the maritime domain which complicate building a maritime SOC (for example, the limited network connectivity of the vessels with shore). The SOC architecture described by the authors took these constraints into account, consisting of vessel-based and shore-based parts. The proposed vessel-based functional blocks of the SOC were the following:
@@
  • Local Preprocessor normalised and correlated events, to lessen the amount of data that needed to be sent to shore via network link via limited bandwidth;
@@
The paper by Jia et al. [55] discussed a big data architecture to collect OT data from Unmanned Surface Vehicles (USVs) for various purposes like business management, disaster prevention, anomaly detection, etc. The authors proposed the collection of the following data from USVs – ship’s performance and navigational data (e.g., speed, position, level of fuel, etc.), data about the environment (e.g., wind, temperature, etc.), data from sensing devices that provided information about the objects around the ship (e.g., data from the radar), and data that was collected from other sources than the USV itself (e.g., industrial data about the USV collected from the manufacturer). For storing the collected data, a cloud-based solution was proposed that was PostgreSQL based. According to the authors, visualisation capabilities were an important aspect of the big data architecture, and the authors suggested the use of the LinDA toolkit for processing the data in the PostgreSQL database. As an example use case for the big data architecture, Jia et al. described an anomaly detection scenario for identifying abnormal combinations of the sailing angle and the USV speed. For anomaly detection purposes, an Isolation Forest-based method was proposed, but the study did not include its performance evaluation.
@@ V-F Security monitoring testbeds
The short paper authored by Becmeur et al. [57] aimed to present a platform to generate data and scenario traces for the evaluation of algorithms for intrusion detection. The introduced solution was based on a ship’s two critical subsystems: 1) the propulsion and engine control in which a computer fan simulated the ship’s propeller and the fan’s actual RPM represented propulsion, 2) the navigation subsystem with the rudder control to simulate the ship’s direction changes. Each subsystem had an independent controller which used common industrial communication protocols, like Modbus, DNP3, and S7. The heart of the setup was the central controller built on a Raspberry Pi. The simulated sensor – that provided the measurement of the engine parts’ temperature – was made of an open-source electronics platform called Arduino. The authors provided an overview of the project, describing its overarching objectives and outcomes, but the overview lacked the required depth for allowing to recreate a similar platform.
@@
The paper by Raimondi et al. [14] described a testbed for conducting cyber exercises to train maritime SOC teams. The testbed followed the maritime SOC architecture proposed by Jacq et al. [13] (see the previous section for a detailed discussion) and was implemented with Linux containerisation techniques. The testbed employed a ship simulator for simulating the ship at sea and a custom Python script for transmitting data received from the simulator as NMEA messages to other ship-related testbed components. For monitoring the ship network, the authors proposed Suricata IDS which employed Lua scripting for parsing NMEA messages. For the cyber situational awareness console on the ship, the authors used Splunk which also forwarded data to a shore-based central SIEM server (another instance of Splunk). To illustrate the training process of maritime SOC operators, the authors described an example cyber exercise which involved the injection of false NMEA messages into the ship network that interfered with the gyrocompass. The task of the trainees was to detect this attack with the help of Splunk query language.
@@
The paper by Sicard et al. [44] described a testbed for experimenting with the Industrial Control System (ICS) of a warship. To achieve a high degree of realism, the authors utilised the physical devices of real ships as much as possible. The testbed implemented the following four areas of the warship – propulsion system (engine and propellers), direction (rudders), energy (fuel for the engine and fuel pump), and artillery (76mm main gun and motors for moving the gun turret). For these four areas, a physical implementation contained a bow and rear of the ship with a physical gun turret, rudders, propellers, physical devices of the bridge, etc. According to the authors, the use of virtualised solutions would have decreased the realism of the testbed, since they would have not allowed for the full imitation of the ship’s real-life environment. For low-level monitoring of Modbus RTU fieldbus traffic in the ICS of the ship, the authors proposed the use of Zeek IDS with specialised traffic capturing hardware and Zeek protocol parser for the Fieldbus protocol (in [69], the authors provided a more detailed description of this solution). For regular network monitoring, commercial IDS sensors were suggested by the authors.
@@
The paper by Longo et al. [38] described the MaCySTe testbed designed for maritime cyber security experiments with the vessel IT/OT systems. The authors released detailed installation instructions of the testbed together with the source code in order to facilitate maritime cyber security research in other organisations. The testbed was built on top of Linux containers, making it a highly resource-efficient solution with modest computational requirements. The testbed provided a detailed emulation of the vessel network, navigational devices, propulsion system, and other components together with relevant protocols (e.g., Navico BR24 and Modbus). For evaluating security monitoring functionality, the testbed supported the use of network probes and SIEM. The network probes captured and parsed traffic in the vessel network, forwarding relevant data to SIEM for visualisation and analysis. The authors also provided some illustrative examples on how the MaCySTe testbed can be utilised for cyber attack and attack detection experiments.
@@ VII Conclusion and future work
As for future work, we plan to study the methods of creating a realistic maritime cyber security monitoring lab, using our experience in the field [27, 28]. Also, we plan to use this lab environment for the creation and release of maritime-specific datasets to facilitate security monitoring research in the maritime domain. Finally, our plans include studying advanced ML-based security monitoring methods on these datasets.

### v3

@@ V-D Security monitoring methods for maritime CPS
In [49], graph-based models were utilised for the detection of attack propagation in CPSs. A graph model that represented the digital and physical system layers was proposed in the paper, considering the interdependencies between CPSs by using the system variables. The model was applied in a case study that involved a small-scale system consisting of a ship with a fuel tank, valve, motor and propeller which were monitored through a PLC. One scenario of normal system usage and four attack scenarios including port scan, closing the PLC and man-in-the-middle attacks were simulated in the study. According to the authors, the method did not produce a false positive alert for the normal scenario, whereas one attack scenario out of four remained undetected (i.e., precision and recall were 100% and 75%, respectively). However, the evaluation of the method was limited to the aforementioned five scenarios without testing the method on a larger dataset.
@@
The study [48] focused on the use of Teager-Kaiser operator for time series analysis to detect anomalies in a ship’s propulsion system in the case of cyber attacks. During the experiments, the authors collected the following time series data – oil consumption, fuel consumption, propeller speed, and valve opening. The following attacks were conducted for evaluating the anomaly detection methods – port scanning with nmap tool, DoS attack with hping tool, and exploiting the vulnerability of the ship’s propulsion PLC for stopping and starting the PLC. Port scanning and DoS attacks did not influence the propulsion system, and only the attack against the PLC had a significant impact. According to the experiments, the Teager-Kaiser operator managed to identify sudden changes in the ship’s propulsion system time series data. However, the evaluation results were presented in the form of graphs only for the 5-minute time period which involved a single attack against the PLC. Therefore, it is difficult to assess what would be the detection rate for many attacks of different types against the propulsion system during a longer time frame (e.g., several days or weeks), and a more detailed analysis of the proposed method was identified as future work in the study.
The study [46] proposed the use of a Particle filter-based anomaly detection method for time series data collected from a ship’s CPS to detect cyber attacks. In the study, the authors focused on data collected from the ship’s propulsion system – the temperature of the cylinder exhaust, cooling water pressure, and the speed of the fuel flow. For training the anomaly detection system, the authors used data collected from a new cargo ship during its first sailing week, assuming that the probability of cyber attacks during this week is very low. For evaluating the anomaly detection algorithm, data from the second half of the first sailing year was employed. For discussing the anomaly detection rate of the algorithm, the authors presented graphs for the 5,000-minute time frame (about 3.5 days) in the study. Although the graphs displayed some fluctuations in the time series data for the propulsion system, the study did not detail whether these sudden changes were the result of a cyber attack or rather a side effect of normal navigation.
@@ V-E Security monitoring system architectures
In [13], a maritime SOC architecture was proposed. The authors first provided an overview of the specific constraints of the maritime domain which complicate building a maritime SOC (for example, the limited network connectivity of the vessels with shore). The SOC architecture described by the authors took these constraints into account, consisting of vessel-based and shore-based parts. The proposed vessel-based functional blocks of the SOC were the following:
@@
  • Local Preprocessor normalised and correlated events, to lessen the amount of data that needed to be sent to shore via network link with limited bandwidth;
@@
The paper [55] discussed a big data architecture to collect OT data from Unmanned Surface Vehicles (USVs) for various purposes like business management, disaster prevention, anomaly detection, etc. The authors proposed the collection of the following data from USVs – ship’s performance and navigational data (e.g., speed, position, level of fuel, etc.), data about the environment (e.g., wind, temperature, etc.), data from sensing devices that provided information about the objects around the ship (e.g., data from the radar), and data that was collected from other sources than the USV itself (e.g., industrial data about the USV collected from the manufacturer). For storing the collected data, a cloud-based solution was proposed that was using PostgreSQL database. According to the authors, visualisation capabilities were an important aspect of the big data architecture, and the authors suggested the use of the LinDA toolkit for processing the data in the PostgreSQL database. As an example use case for the big data architecture, the paper described an anomaly detection scenario for identifying abnormal combinations of the sailing angle and the USV speed. For anomaly detection purposes, an Isolation Forest-based method was proposed, but the study did not include its performance evaluation.
@@ V-F Security monitoring testbeds
The short paper [57] aimed to present a platform to generate data and scenario traces for the evaluation of algorithms for intrusion detection. The introduced solution was based on a ship’s two critical subsystems: 1) the propulsion and engine control in which a computer fan simulated the ship’s propeller and the fan’s actual RPM represented propulsion, 2) the navigation subsystem with the rudder control to simulate the ship’s direction changes. Each subsystem had an independent controller which used common industrial communication protocols, like Modbus, DNP3, and S7. The heart of the setup was the central controller built on a Raspberry Pi. The simulated sensor – that provided the measurement of the engine parts’ temperature – was made of an open-source electronics platform called Arduino. The authors provided an overview of the project, describing its overarching objectives and outcomes, but the overview lacked the required depth for allowing to recreate a similar platform.
@@
Raimondi et al. [14] described a testbed for conducting cyber exercises to train maritime SOC teams. The testbed followed the maritime SOC architecture proposed in [13] (see the previous section for a detailed discussion) and was implemented with Linux containerisation techniques. The testbed employed a ship simulator for simulating the ship at sea and a custom Python script for transmitting data received from the simulator as NMEA messages to other ship-related testbed components. For monitoring the ship network, the authors proposed Suricata IDS which employed Lua scripting for parsing NMEA messages. For the cyber situational awareness console on the ship, the authors used Splunk which also forwarded data to a shore-based central SIEM server (another instance of Splunk). To illustrate the training process of maritime SOC operators, the authors described an example cyber exercise which involved the injection of false NMEA messages into the ship network that interfered with the gyrocompass. The task of the trainees was to detect this attack with the help of Splunk query language.
@@
In [44], a testbed was presented for experimenting with the Industrial Control System (ICS) of a warship. To achieve a high degree of realism, the authors utilised the physical devices of real ships as much as possible. The testbed implemented the following four areas of the warship – propulsion system (engine and propellers), direction (rudders), energy (fuel for the engine and fuel pump), and artillery (76mm main gun and motors for moving the gun turret). For these four areas, a physical implementation contained a bow and rear of the ship with a physical gun turret, rudders, propellers, physical devices of the bridge, etc. According to the authors, the use of virtualised solutions would have decreased the realism of the testbed, since they would have not allowed for the full imitation of the ship’s real-life environment. For low-level monitoring of Modbus RTU fieldbus traffic in the ICS of the ship, the authors proposed the use of Zeek IDS with specialised traffic capturing hardware and Zeek protocol parser for the Fieldbus protocol (in [69], the authors provided a more detailed description of this solution). For regular network monitoring, commercial IDS sensors were suggested by the authors.
@@
Longo et al. [38] described the MaCySTe testbed designed for maritime cyber security experiments with the vessel IT/OT systems. The authors released detailed installation instructions of the testbed together with the source code in order to facilitate maritime cyber security research in other organisations. The testbed was built on top of Linux containers, making it a highly resource-efficient solution with modest computational requirements. The testbed provided a detailed emulation of the vessel network, navigational devices, propulsion system, and other components together with relevant protocols (e.g., Navico BR24 and Modbus). For evaluating security monitoring functionality, the testbed supported the use of network probes and SIEM. The network probes captured and parsed traffic in the vessel network, forwarding relevant data to SIEM for visualisation and analysis. The authors also provided some illustrative examples on how the MaCySTe testbed can be utilised for cyber attack and attack detection experiments.
@@ VII Conclusion and future work
As for future work, we plan to study the methods of creating a realistic maritime cyber security monitoring lab, using our experience in the field [27]. Also, we plan to use this lab environment for the creation and release of maritime-specific datasets to facilitate security monitoring research in the maritime domain. Finally, our plans include studying advanced ML-based security monitoring methods on these datasets.